It is January 9, 2017. The price of a Treasury bond with a 6% coupon that matures on Oct 12, 2030, is quoted as 102-04. What is the cash price?

(This bond pays coupons on Oct 12 and Apr 12. Excel might be helpful in calculating the difference in dates.)

(required precision: 0.01 +/- 0.01)
Answer #1

Number of days between october 12, 2014 and January 9, 2015 = 89 days

Number of days between october 12, 2014 and April 12, 2015 = 182 days

Quoted price = 102 + 04 / 32 = 102.125

Accrued interest = 100 * 6% * (6/12) * (89/182) = 1.467

Cash price of bond = Accrued Interest + Quoted price
= 1.467 + 102.125
= 103.59

Cash price of bond = 103.59